OpenStudy (hihi):

Last question: What is the slope of the line passing through the point (3, −4) and perpendicular to the line y = 3x + 5?

OpenStudy (anonymous):

homework almost done?

OpenStudy (anonymous):

The slope of your line is 3/1 Do you know the slope of the perpendicular line?

OpenStudy (hihi):

Well im almost done, ive got 5 problems left, but il do them by myself, this question is hard for me though.

OpenStudy (hihi):

the slope of the perpendicular line is 1/3 right?

OpenStudy (hihi):

because isnt it supposed to be the opposite?

OpenStudy (anonymous):

-1/3

OpenStudy (karatechopper):

no the perpendicular is the opposite sign of the reciprocal

OpenStudy (hihi):

oh ya, forgot to add the negative, but i remembered it.

OpenStudy (anonymous):

Yes, the perpendicular slope is -1/3. We will use the point-slope formula y - y1 = m(x - x1) we know the slope is -1/3 and we have x = 3 and y = -4; therefore, y + 4 = (-1/3)(x - 3) Is this form oaky?

OpenStudy (hihi):

THX FOR ALL YOUR HELP TODAY GUYS

OpenStudy (hihi):

especially sam unleashed, meverett and karatechoppers

OpenStudy (anonymous):

u welcome

OpenStudy (hihi):

thx